A factory worker places 120 newly created circuits on a shelf to be checked for quality. Of these, 9 will not work correctly. Suppose that she is asked to randomly select two circuits, without replacement, from the shelf. What is the chance that both circuits she selects will be defective? Show step by step work! Approximate the solution to the nearest ten thousandth.



Answer by checkley77(12844)

9/120*8/119=72/14,280=.0053 or .53% chance of selecting 2 defective circuits.